.A man whose eyes are 130 cm above the oil film on water surface observes greenish colour at a distance of 90 c.m from his feet. Calculate the probable thickness of the film. (green=5000A0 ,µoil=1.4, µwater=1.33).


1
Expert's answer
2021-07-05T08:43:24-0400

"\frac {\lambda}{\mu_w d}=\frac x{\mu_o l},"

"x=\lambda\frac{\mu_o}{\mu_w}\frac ld,"

"x=5\cdot 10^{-7}\cdot \frac{1.4}{1.33}\cdot \frac{130}{90}=0.76~\mu m."